OpenStudy (anonymous):

The absolute value equation |4(x – 1)| = 12 has two solutions. True False

OpenStudy (anonymous):

false

OpenStudy (anonymous):

true

OpenStudy (lgbasallote):

when you have \[| \; f(x) \; | = a\] then there are automatically two solutions... however.. if you have \[| \; f(x) \; | = -a\] then there are no solutions does that help?

OpenStudy (anonymous):

yep... |4(x-1)|=12 4(x-1)=-12 x<0 4(x-1)=12 x>0 x-1=-3 or x-1=3 x=-2 or 4

OpenStudy (anonymous):

so its false :)

OpenStudy (lgbasallote):

it's true...

OpenStudy (anonymous):

\[\log_{}\left| x \right| =2\] how many solutions

OpenStudy (lgbasallote):

all three of us said true and you decided to go with the answer of the only one who said false lol

OpenStudy (anonymous):

oh gotcha :) sorry bout that :)

OpenStudy (anonymous):

it has only a single soln since x cannot be -ve in case of log

OpenStudy (anonymous):

@Jonask

OpenStudy (lgbasallote):

it can because of the absolute value

OpenStudy (anonymous):

I'm blonde lol

OpenStudy (lgbasallote):

that's why they put absolute value on logs...to denote that the argument can't be negative

OpenStudy (anonymous):

oh so x can have two hear but at the it becomes +

OpenStudy (lgbasallote):

yes it does

OpenStudy (anonymous):

thanks